OpenStudy (anonymous):

Find the lateral area of a cone with altitude (height) 6 cm and base diameter 16 cm.

OpenStudy (anonymous):

LSA= 1/2 (l)(2)pi(r) =1/2 (6)(2)pi(8) =48pi

OpenStudy (anonymous):

the possible answers are 80 pi 120 pi 144 pi 160 pi